# A framework for fluctuating times and counting observables in stochastic excursions

### v1

@@ Appendix B Further comments on the absorption refrigerator
Figure 7: (a) Cold current as a function of cold reservoir occupation number $n_{c}$, for the classical and quantum dynamics. We considered $n_{w}=0.62$ fixed. (b) Population of the cold qubit over time for different values of $n_{w}$, with fixed $n_{c}=0.028$. Solid lines are obtained from the classical master equation and dashed lines from the quantum master equation. All parameters were chosen to match exactly the ones in Ref. aamirThermallyDrivenQuantum2025 . We fixed, in units of MHz/$2\pi$, the parameters: $\Gamma_{h}=7.2,\leavevmode\nobreak\ \Gamma_{w}=0.07,\leavevmode\nobreak\ \Gamma_{c}=0.0087,\leavevmode\nobreak\ g=3.2,\leavevmode\nobreak\ \omega_{w}=5.327\times 10^{3},\leavevmode\nobreak\ \omega_{c}=3.725\times 10^{3},\leavevmode\nobreak\ \omega_{h}=\omega_{w}+\omega_{c}$. We fixed the occupation number of the hot reservoir in both plots to be $n_{h}=0.007$.
@@
$J_{\rm quant}=2g\omega_{\alpha}\leavevmode\nobreak\ {\rm Im}\left[\mathrm{tr}{(\sigma_{c}^{+}\sigma_{h}^{-}\sigma_{h}^{+}\rho^{\rm ss}})\right],$ (128)
@@
Clearly, the results for the current $J_{c}$ and the real-time dynamics of the population of the cold qubit $p_{c}$ suggest that indeed the classical master equation (1) is a very good approximation of the full quantum dynamics for the three qubit absorption refrigerator. We emphasize that the parameters used in the plots of Fig. 7 were chosen to match the experimental values of Ref. aamirThermallyDrivenQuantum2025 . Of course, the parameter regime used in the simplified model of Ref. Fiusa2025 and the full model in this paper [see results in Figs. 3 and 4] also work very well in the classical approximation.

### v2

@@ Appendix B Further comments on the absorption refrigerator
[…]
@@
$J_{\rm quant}=2g\omega_{\alpha}\penalty 10000\ {\rm Im}\left[\mathrm{tr}{(\sigma_{c}^{+}\sigma_{h}^{-}\sigma_{h}^{+}\rho^{\rm ss}})\right],$ (139)
@@
Clearly, the results for the current $J_{c}$ and the real-time dynamics of the population of the cold qubit $p_{c}$ suggest that indeed the classical master equation (1) is a very good approximation of the full quantum dynamics for the three qubit absorption refrigerator. We emphasize that the parameters used in the plots of Fig. 8 were chosen to match the experimental values of Ref. aamirThermallyDrivenQuantum2025 . Of course, the parameter regime used in the simplified model of Ref. Fiusa2025 and the full model in this paper [see results in Figs. 4 and 5] also work very well in the classical approximation.